Calculate how much of each solution (and water!) you would need to make 100 mLs of a 0.1 M Tris/0.05M EDTA solution from 1M Tris and 1 M EDTA stock solutions?
Answer #1

Using the law of conservation of mass,

M1. V1 = M2 .V2

(1 denotes the values from the dilution we require and 2 denotes the values of stock solution)

  • 100mL of 0.1M tris from 1M tris stock

0.1 * 100 = 1 * x

therefore x = 0.1 * 100/ 1

= 10 mL

therefore take 10mL from tris stock and mix it with 90mL water

  • 100mL of 0.05 M EDTA from 1M EDTA stock

0.05 * 100 = 1 * x

x = 0.05 * 100/1

x= 5mL

thereforetake 5mL from EDTA stock and mix with 95mL water to get 100mL of 0.05M EDTA